Costs per Equivalent Unit and Production Costs

The following information concerns production in the Forging Department for November. All direct materials are placed into the process at the beginning of production, and conversion costs are incurred evenly throughout the process. The beginning inventory consists of $123,200 of direct materials.

ACCOUNT Work in Process—Forging Department ACCOUNT NO.
Date Item Debit Credit Balance
Debit Credit
Nov. 1 Bal., 8,800 units, 30% completed 132,704
30 Direct materials, 80,000 units 1,104,000 1,236,704
30 Direct labor 129,540 1,366,244
30 Factory overhead 178,892 ? 1,545,136
30 Goods finished, ? units ?
30 Bal., 7,000 units, 60% completed ?

Cost per equivalent units of $13.80 for Direct Materials and $3.70 for Conversion Costs.

Based on the above data, determine each of the following amounts.

If required, round your interim calculations to two decimal places. Round final answers (a-c) to the nearest dollar.

a. Cost of beginning work in process inventory completed in November.
$

b. Cost of units transferred to the next department during November.
$

c. Cost of ending work in process inventory on November 30.
$

d. Costs per equivalent unit of direct materials and conversion included in the November 1 beginning work in process. If required, round your answers to two decimal places.

Direct materials cost per equivalent unit $
Conversion cost per equivalent unit $

e. The November increase or decrease in costs per equivalent unit for direct materials and conversion from the previous month. If required, round your answers to two decimal places.

Increase or Decrease Amount
Change in direct materials cost per equivalent unit $
Change in conversion cost per equivalent unit $
Answer #1

to solve this question we need to calculate some figures which are showing in the form of statement showing below.

statement of equivalent production units

particular input units particular output units equivalent production

% units

opening wip 8800 from opening wip 8800 70% 6160

units introduced 80,000 from fresh inputs 80,000   

units completed and

transfer 73000 100 73000

closing wip 7000 60 4200

total 88800 88800 83360

answer d. statement showing cost for each element

particular material conversion cost total

cost of opening wip 123200 132704 255904

cost incurred during month 1104000 308432 1412432

total cost 1227200 441136   

equivalent units 83360 83360

cost per equivalent units 14.72 5.29

statement of apportionment of cost

answer a. value of opening wip 123200+132704=255904

answer b. value of units introduced and transfer   

material 73000*14.72 1074560

conversion cost 73000*5.29 386170

total cost of units completed and transfer 255904+1074560+386170=1716634

answer c. cost of closing wip  

material 7000*14.72 = 103040

conversion cost 7000*5.29 = 37030

total 140070

Answer e. change in equivalent cost per unit previous present change

change in direct material equivalent cost per unit 13.80 14.72 0.9

change in conversion cost 3.70 5.29 1.59